OpenStudy (anonymous):

A car rental agency charges $175 per week plus $0.20 per mile to rent a car. How many miles can you travel in one week for A) 238 miles B) 675 miles C) 700 miles D) 1575 miles

OpenStudy (anonymous):

$315

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

ur a pain in my head

OpenStudy (anonymous):

for the previous problem we had p = 0.40*(x-1)+2.30 can u change it to ur problem

OpenStudy (anonymous):

ok. lets see.

OpenStudy (anonymous):

0.40 was the price for each paper 2.30 was for the 1st one

OpenStudy (anonymous):

350=.20(x-1)-175

OpenStudy (anonymous):

thats suppose to be 315

OpenStudy (anonymous):

490=.20x-.2

OpenStudy (anonymous):

ok ...

OpenStudy (anonymous):

its 350=.20(x-1)+175

OpenStudy (anonymous):

in fact we dont need the one because the 175 doesnt include the 1st mile so 315 = 0.20(x) + 175

OpenStudy (anonymous):

315 = 0.20(x) + 175 -175 -175 140 = 0.20x

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

7

OpenStudy (anonymous):

imm lost!

OpenStudy (anonymous):

what do u divide with?

OpenStudy (anonymous):

140/.20

OpenStudy (anonymous):

use calculator

OpenStudy (anonymous):

oh 700